OpenStudy (anonymous):

How to prove that a/b 0 means that a/b<(a+c)/(b+d)

OpenStudy (phi):

Some background: Given all numbers positive and\[ \frac{a}{b} < \frac{c}{d}\] if we multiply both sides by b, the relation remains true. ditto for d. thus \[ ad < bc\] It is not obvious how to use this observation, but let's work backwards. The claim is \[ \frac{a}{b} < \frac{a+c}{b+d} \] as before "cross multiply" i.e. multiply both sides by the denominators: \[ ab+ad < ab +bc \] subtracting ab from both sides, and we have \[ ad <bc\] which implies \( \frac{a}{b} < \frac{c}{d}\) so we know how to proceed: \[ \frac{a}{b} < \frac{c}{d} \implies ad<bc \implies ab+ad< ab+bc \\ \implies a(b+d) < b(a+c) \implies \frac{a}{b} < \frac{a+c}{b+d} \] we can do a similar thing for the upper bound

OpenStudy (anonymous):

Given: a/b < c/d; a,b,c,d>0 We know that if a/b = c/d then ad = bc; =>a/b = (a+c)/(b+d). when you cross multiply you get the same result ad = bc. This hold true for the inequality since a,b,c,d>0 That is a/b<c/d => a/b < (a+c)/(b+d) .......(1) Now, we want to see whether (a+c)/(b+d) < c/d Let us assume (a+c)/(b+d)>c/d. When you cross multiply you get ad>bc, which is contradictory to our initial assumption. Therefore, our assumption is wrong. It implies (a+c)/(b+d)<c/d ........(2) Combining (1) and (2), we get a/b<(a+c)/(b+d)<c/d Hence, proved.